# A Progressive Image Restoration Network for High-order Degradation Imaging in Remote Sensing

### v1

@@ IV-B5 DDLB architecture
TABLE VI: Comparison of different SRx2 models on various scene classes of WHU-RS19 (PSNR ($dB$)/SSIM). The best and second-best results are highlighted in Bold font and underlined ones, respectively.
WHU-RS19 Airport Beach Bridge Commercial Desert Farmland FootballField Forest Industrial Meadow
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 27.08/0.737 34.41/0.944 31.79/0.854 24.31/0.693 36.79/0.898 33.52/0.824 28.05/0.793 27.03/0.586 25.54/0.718 34.81/0.815
HSENet [49] 27.55/0.745 34.41/0.950 32.25/0.858 24.60/0.694 36.83/0.901 33.95/0.828 28.39/0.799 27.18/0.594 25.94/0.723 34.84/0.813
D-DBPN [45] 27.80/0.758 35.67/0.945 32.54/0.863 24.74/0.704 37.03/0.900 33.82/0.828 28.76/0.809 27.39/0.607 26.12/0.734 34.95/0.815
RDN [17] 27.90/0.761 36.18/0.948 32.65/0.867 24.79/0.706 37.29/0.903 34.16/0.835 28.89/0.818 27.44/0.612 26.22/0.739 35.06/0.817
TransENet [28] 27.92/0.762 36.28/0.946 32.54/0.867 24.84/0.708 37.33/0.903 33.99/0.832 29.09/0.820 27.47/0.617 26.37/0.745 34.96/0.814
RCAN [10] 27.97/0.765 36.73/0.949 32.83/0.870 24.94/0.714 37.43/0.904 34.28/0.838 29.14/0.823 27.41/0.612 26.41/0.747 35.15/0.820
SAN [20] 27.98/0.761 36.64/0.943 32.49/0.861 24.87/0.709 37.22/0.903 34.27/0.835 28.97/0.817 27.45/0.612 26.29/0.740 35.12/0.819
MHAN [24] 27.91/0.762 36.49/0.951 32.58/0.865 24.88/0.711 37.38/0.904 34.23/0.835 28.96/0.820 27.43/0.613 26.37/0.743 35.02/0.816
SwinIR [12] 27.96/0.764 37.18/0.950 32.62/0.868 24.88/0.712 37.44/0.904 34.24/0.836 29.00/0.823 27.36/0.612 26.33/0.747 35.13/0.819
HAUNet [27] 28.05/0.762 36.60/0.948 32.81/0.870 24.97/0.711 37.50/0.904 34.40/0.838 29.24/0.823 27.50/0.619 26.53/0.747 35.17/0.820
TTST [26] 27.97/0.763 36.59/0.944 32.72/0.869 24.93/0.712 37.52/0.905 34.26/0.835 29.05/0.821 27.39/0.611 26.38/0.746 35.11/0.819
CAT [46] 28.10/0.768 37.15/0.954 32.93/0.871 25.11/0.720 37.53/0.905 34.36/0.838 29.42/0.832 27.57/0.626 26.56/0.755 35.15/0.819
RGT [47] 28.19/0.771 37.18/0.949 32.83/0.872 25.09/0.719 37.63/0.906 34.50/0.841 29.37/0.831 27.50/0.624 26.56/0.754 35.21/0.821
DAT [22] 28.25/0.772 36.99/0.948 32.83/0.873 25.11/0.721 37.65/0.906 34.53/0.841 29.48/0.832 27.52/0.624 26.60/0.756 35.20/0.821
HDI-PRNet 28.55/0.780 37.80/0.950 33.48/0.878 25.44/0.731 37.58/0.906 34.59/0.845 29.73/0.836 27.84/0.634 27.08/0.767 35.22/0.822
WHU-RS19 Mountain Park Parking Pond Port RailwayStation Residential River Viaduct Average
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 24.46/0.579 27.44/0.701 25.81/0.805 35.10/0.902 26.62/0.806 24.56/0.621 24.20/0.724 27.63/0.699 25.17/0.693 28.64/0.757
HSENet [49] 24.66/0.582 27.84/0.704 26.52/0.817 35.57/0.905 27.03/0.811 24.81/0.623 24.48/0.730 27.77/0.703 25.45/0.701 28.94/0.762
D-DBPN [45] 24.81/0.592 27.91/0.708 26.92/0.830 35.56/0.901 27.31/0.821 25.12/0.644 24.74/0.742 27.94/0.710 25.96/0.725 29.20/0.770
RDN [17] 24.83/0.594 27.99/0.711 26.93/0.833 35.82/0.907 27.31/0.824 25.19/0.653 24.74/0.745 28.01/0.714 26.04/0.732 29.33/0.774
TransENet [28] 24.87/0.604 28.03/0.713 26.94/0.836 35.65/0.907 27.33/0.825 25.08/0.646 24.82/0.748 27.98/0.719 26.09/0.733 29.33/0.776
RCAN [10] 24.92/0.602 28.09/0.716 27.01/0.838 35.97/0.909 27.39/0.827 25.21/0.658 24.90/0.752 28.02/0.719 26.24/0.740 29.46/0.779
SAN [20] 24.87/0.599 28.05/0.712 26.92/0.833 35.82/0.907 27.35/0.825 25.17/0.650 24.83/0.747 28.09/0.716 26.03/0.729 29.38/0.774
MHAN [24] 24.89/0.600 28.07/0.713 27.14/0.837 35.81/0.906 27.48/0.828 25.19/0.650 24.89/0.750 28.10/0.718 26.13/0.733 29.40/0.776
SwinIR [12] 24.88/0.604 28.05/0.716 26.94/0.839 35.91/0.907 27.38/0.829 25.01/0.652 24.87/0.751 27.94/0.718 26.13/0.738 29.42/0.778
HAUNet [27] 25.00/0.607 28.10/0.715 27.31/0.844 36.03/0.910 27.43/0.828 25.29/0.656 24.90/0.751 28.09/0.720 26.25/0.736 29.52/0.779
TTST [26] 24.86/0.600 28.07/0.714 27.13/0.840 35.91/0.909 27.43/0.829 25.31/0.651 24.92/0.752 28.07/0.719 26.09/0.733 29.44/0.777
CAT [46] 24.95/0.608 28.18/0.721 27.46/0.850 35.99/0.909 27.70/0.834 25.38/0.661 25.10/0.761 28.18/0.725 26.46/0.746 29.63/0.784
RGT [47] 24.95/0.610 28.25/0.722 27.27/0.845 36.02/0.910 27.67/0.834 25.31/0.662 25.10/0.761 28.15/0.725 26.46/0.749 29.63/0.784
DAT [22] 25.08/0.615 28.25/0.722 27.39/0.847 36.16/0.911 27.76/0.836 25.50/0.668 25.16/0.763 28.22/0.728 26.53/0.752 29.68/0.785
HDI-PRNet 25.27/0.627 28.48/0.727 28.18/0.860 36.24/0.911 28.35/0.845 26.05/0.693 25.58/0.772 28.54/0.735 26.87/0.761 30.03/0.793
@@
TABLE VII: Comparison of different SRx3 models on various scene classes of WHU-RS19 (PSNR ($dB$)/SSIM). The best and second-best results are highlighted in Bold font and underlined ones, respectively.
WHU-RS19 Airport Beach Bridge Commercial Desert Farmland FootballField Forest Industrial Meadow
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 26.10/0.716 34.47/0.945 29.89/0.843 23.08/0.636 36.60/0.895 32.52/0.800 26.78/0.757 26.64/0.556 24.47/0.665 34.10/0.802
HSENet [49] 26.59/0.726 35.66/0.942 30.47/0.851 23.36/0.645 36.78/0.896 33.00/0.808 27.22/0.769 26.75/0.562 24.89/0.679 34.18/0.803
D-DBPN [45] 26.82/0.739 35.57/0.948 31.01/0.860 23.96/0.673 36.76/0.894 32.92/0.809 27.77/0.785 26.95/0.580 25.44/0.702 34.40/0.807
RDN [17] 26.90/0.741 36.40/0.951 31.07/0.861 23.89/0.671 37.01/0.899 33.26/0.813 27.73/0.790 26.93/0.578 25.39/0.701 34.42/0.807
TransENet [28] 27.09/0.744 36.91/0.953 31.06/0.864 23.94/0.674 36.86/0.899 33.18/0.813 27.79/0.790 27.01/0.587 25.49/0.704 34.34/0.806
RCAN [10] 27.09/0.748 36.57/0.948 31.18/0.865 23.99/0.680 37.03/0.899 33.29/0.814 27.87/0.794 27.07/0.593 25.49/0.707 34.53/0.810
SAN [20] 26.86/0.741 35.83/0.944 30.90/0.861 23.88/0.671 36.95/0.898 33.20/0.812 27.68/0.785 26.91/0.578 25.41/0.701 34.44/0.807
MHAN [24] 27.01/0.745 36.20/0.951 31.18/0.864 23.99/0.676 36.99/0.898 33.26/0.814 27.84/0.791 26.99/0.585 25.53/0.706 34.46/0.809
SwinIR [12] 27.10/0.749 37.11/0.954 31.21/0.866 24.00/0.680 37.08/0.900 33.30/0.814 27.81/0.795 27.06/0.592 25.53/0.709 34.48/0.809
HAUNet [27] 27.09/0.745 36.61/0.955 31.21/0.864 23.98/0.673 37.02/0.899 33.38/0.815 27.88/0.788 27.09/0.588 25.56/0.704 34.46/0.807
TTST [26] 27.00/0.743 35.63/0.955 31.13/0.862 23.92/0.671 36.69/0.898 33.31/0.814 27.85/0.787 27.00/0.582 25.47/0.701 34.46/0.808
CAT [46] 27.35/0.755 36.94/0.949 31.37/0.868 24.16/0.687 37.13/0.900 33.45/0.818 28.16/0.802 27.13/0.595 25.69/0.715 34.58/0.810
RGT [47] 27.31/0.755 37.29/0.955 31.37/0.869 24.16/0.687 37.14/0.900 33.47/0.819 28.19/0.804 27.13/0.596 25.69/0.716 34.54/0.809
DAT [22] 27.34/0.756 37.07/0.953 31.53/0.870 24.20/0.689 37.15/0.900 33.51/0.820 28.26/0.804 27.14/0.598 25.75/0.718 34.56/0.810
HDI-PRNet 27.53/0.759 37.62/0.957 32.02/0.874 24.41/0.692 36.97/0.900 33.66/0.822 28.61/0.808 27.23/0.596 26.02/0.723 34.60/0.811
WHU-RS19 Mountain Park Parking Pond Port RailwayStation Residential River Viaduct Average
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 23.53/0.501 26.72/0.653 24.04/0.735 32.81/0.883 25.21/0.785 23.87/0.575 22.79/0.660 26.57/0.648 24.22/0.638 27.59/0.720
HSENet [49] 23.74/0.514 27.05/0.661 24.59/0.749 33.12/0.888 25.48/0.792 24.31/0.596 23.17/0.673 26.93/0.658 24.59/0.653 27.98/0.729
D-DBPN [45] 23.93/0.525 27.31/0.672 25.27/0.775 33.36/0.890 26.06/0.805 24.78/0.623 23.57/0.693 27.14/0.668 25.10/0.682 28.31/0.743
RDN [17] 23.95/0.533 27.35/0.673 25.19/0.776 33.46/0.893 25.97/0.806 24.80/0.626 23.46/0.692 27.19/0.669 25.08/0.684 28.38/0.745
TransENet [28] 23.97/0.540 27.38/0.675 25.54/0.788 33.32/0.893 25.98/0.808 24.91/0.628 23.63/0.700 27.23/0.675 25.09/0.686 28.44/0.748
RCAN [10] 24.03/0.544 27.46/0.679 25.17/0.784 33.55/0.895 26.03/0.812 24.96/0.640 23.51/0.699 27.30/0.678 25.28/0.696 28.48/0.751
SAN [20] 23.91/0.530 27.35/0.673 25.12/0.774 33.41/0.892 25.89/0.805 24.73/0.623 23.50/0.694 27.14/0.668 25.03/0.679 28.31/0.743
MHAN [24] 24.00/0.538 27.43/0.676 25.23/0.782 33.57/0.894 26.09/0.810 24.94/0.634 23.58/0.698 27.25/0.673 25.21/0.688 28.45/0.748
SwinIR [12] 24.05/0.547 27.44/0.679 25.44/0.790 33.55/0.895 26.14/0.814 24.97/0.637 23.61/0.703 27.29/0.677 25.26/0.696 28.53/0.752
HAUNet [27] 24.04/0.539 27.35/0.671 25.64/0.792 33.54/0.893 26.07/0.808 24.93/0.627 23.65/0.700 27.33/0.675 25.16/0.685 28.51/0.748
TTST [26] 23.98/0.535 27.40/0.674 25.39/0.779 33.41/0.892 26.04/0.807 24.77/0.621 23.60/0.698 27.26/0.674 25.17/0.686 28.38/0.746
CAT [46] 24.13/0.553 27.47/0.681 25.80/0.804 33.59/0.896 26.33/0.818 25.18/0.647 23.77/0.710 27.38/0.683 25.47/0.707 28.67/0.757
RGT [47] 24.09/0.550 27.57/0.683 25.66/0.799 33.73/0.897 26.34/0.819 25.25/0.652 23.77/0.710 27.41/0.684 25.47/0.708 28.70/0.758
DAT [22] 24.12/0.552 27.53/0.682 25.89/0.803 33.71/0.897 26.37/0.819 25.18/0.651 23.79/0.712 27.44/0.684 25.52/0.709 28.72/0.759
HDI-PRNet 24.24/0.550 27.65/0.684 26.59/0.818 33.74/0.896 26.82/0.825 25.55/0.662 24.31/0.722 27.53/0.685 25.66/0.711 28.96/0.762

### v2

@@ IV-B5 DDLB architecture
TABLE VII: Comparison of different SR $\times$2 models on various scene classes of WHU-RS19 (PSNR ($dB$)). The best and second-best results are highlighted in Bold font and underlined ones, respectively.
Class CTNet HSENet D-DBPN RDN TransE- RCAN SAN MHAN SwinIR HAUNet TTST CAT RGT DAT HDI-
[58] [59] [55] [22] Net[33] [14] [25] [29] [17] [32] [31] [56] [57] [27] PRNet
Airport 27.08 27.55 27.80 27.90 27.92 27.97 27.98 27.91 27.95 28.05 27.97 28.10 28.19 28.25 28.55
Beach 34.41 34.40 35.67 36.18 36.28 36.73 36.64 36.49 37.11 36.60 36.55 37.15 37.18 36.99 37.80
Bridge 31.79 32.25 32.54 32.65 32.54 32.83 32.49 32.58 32.58 32.81 32.83 32.93 32.83 32.83 33.48
Commercial 24.31 24.60 24.74 24.79 24.84 24.94 24.87 24.88 24.88 24.97 24.94 25.11 25.09 25.11 25.44
Desert 36.79 36.83 37.03 37.29 37.33 37.43 37.22 37.38 37.46 37.50 37.38 37.53 37.63 37.65 37.58
Farmland 33.52 33.95 33.82 34.16 33.99 34.28 34.27 34.23 34.23 34.40 34.33 34.36 34.50 34.53 34.59
Football Field 28.05 28.39 28.76 28.89 29.09 29.14 28.97 28.96 29.00 29.24 29.01 29.42 29.37 29.48 29.73
Forest 27.03 27.18 27.39 27.44 27.47 27.41 27.45 27.43 27.36 27.50 27.44 27.57 27.50 27.52 27.84
Industrial 25.54 25.94 26.12 26.22 26.37 26.41 26.29 26.37 26.33 26.53 26.36 26.56 26.56 26.60 27.08
Meadow 34.81 34.84 34.95 35.06 34.96 35.15 35.12 35.02 35.13 35.17 35.13 35.15 35.21 35.20 35.22
Mountain 24.46 24.66 24.81 24.83 24.87 24.92 24.87 24.89 24.88 25.00 24.89 24.95 24.95 25.08 25.27
Park 27.44 27.84 27.91 27.99 28.03 28.09 28.05 28.07 28.04 28.10 28.05 28.18 28.25 28.25 28.48
Parking 25.81 26.52 26.92 26.93 26.94 27.01 26.92 27.14 26.94 27.31 27.18 27.46 27.27 27.39 28.18
Pond 35.10 35.57 35.56 35.82 35.65 35.97 35.82 35.81 35.89 36.03 35.94 35.99 36.02 36.16 36.24
Port 26.62 27.03 27.31 27.31 27.33 27.39 27.35 27.48 27.40 27.43 27.48 27.70 27.67 27.76 28.35
Railway Station 24.56 24.81 25.12 25.19 25.08 25.21 25.17 25.19 25.02 25.29 25.29 25.38 25.31 25.50 26.05
Residential 24.20 24.48 24.74 24.74 24.82 24.90 24.83 24.89 24.86 24.90 24.93 25.10 25.10 25.16 25.58
River 27.63 27.77 27.94 28.01 27.98 28.02 28.09 28.10 27.94 28.09 28.12 28.18 28.15 28.22 28.54
Viaduct 25.17 25.45 25.96 26.04 26.09 26.24 26.03 26.13 26.12 26.25 26.10 26.46 26.46 26.53 26.87
@@
TABLE VIII: Comparison of different SR $\times$2 models on various scene classes of RSSCN7 (PSNR ($dB$)). The best and second-best results are highlighted in Bold font and underlined ones, respectively.
Class CTNet HSENet D-DBPN RDN TransE- RCAN SAN MHAN SwinIR HAUNet TTST CAT RGT DAT HDI-
[58] [59] [55] [22] Net[33] [14] [25] [29] [17] [32] [31] [56] [57] [27] PRNet
Grass 31.81 32.12 32.13 32.27 32.24 32.37 32.31 32.28 32.32 32.37 32.36 32.40 32.46 32.48 32.55
Field 31.38 31.63 31.57 31.75 31.74 31.86 31.82 31.78 31.82 31.90 31.87 31.91 31.98 32.01 32.09
Industry 24.41 24.85 25.03 25.08 25.10 25.17 25.16 25.13 25.11 25.24 25.17 25.28 25.32 25.37 25.75
River Lake 29.24 29.60 29.74 29.79 29.79 29.86 29.83 29.84 29.84 29.86 29.85 29.94 29.97 29.99 30.17
Forest 27.02 27.24 27.31 27.36 27.38 27.43 27.39 27.38 27.41 27.43 27.42 27.49 27.50 27.52 27.61
Resident 23.67 24.06 24.27 24.28 24.35 24.41 24.35 24.35 24.36 24.47 24.37 24.53 24.51 24.54 24.87
Parking 24.18 24.44 24.61 24.62 24.65 24.72 24.69 24.68 24.67 24.76 24.73 24.84 24.87 24.89 25.17